16 November 2007

RHIC Startup:

Vacuum problem in L05.  Beam is not getting through, and vacuum pumps tripped off.  J10 bump not operating for the deuteron cycle.  The puts the beam loss on the inside of the ring.  This damaged the kapton (burned) and broke the vacuum.  It was sprayed with sealant, which allowed the vacuum pumps to be turned on again.  The J10 bump running full current is not enough to keep from losses.

Carl finished his PS work.

Freddy put in new RF scaling factor for the RF code. Blue/Yellow synchronization has been fixed.

Vadim found a vertical corrector with the wrong polarity. It has been put on STBY and will be fixed today.  It may be responsible for the trouble making orbit corrections.  He also found a bad BPM and put it off.

Tune/Coupling feedback loops need to be closed. Closing the loops now appear to increase the coupling rather than decrease it.  So until Yun can work on the tune/coupling feedback, we will work just on improving the ramp and tune.  BBQ locks fine on Blue.  Getting BBQ to work on Yellow would be good.

 


Scheduling Physicist: Mike Sivertz Last updated:
Email:sivertz@bnl.gov
Office Phone: 631-344-6102
Digital Pager: 7528 or 4300
Home Phone: 631-878-9020